Minutes of the Tracker Meeting

November 11, 2004


      Action Items:

1.    Dave: arrange a date for the Teledyne delta-MRR.

2.    Dave: get Zentek’s payment squared away so that we can move forward with cable rework.

3.    Jeff: check on the Poron foam order.

4.    Jeff: send a non-flight cable to Pisa for fit checks to the top tray hardware.

5.    Jeff: find or purchase the other halves of the long GASU cables.

6.    Jeff and INFN: find the missing long PDU cables.

7.    All: review and send approvals for the tray assembly drawings.

8.    Sandro, Riccardo, and Ben: review and approve PS-03467 (sidewall specs).

9.    Jeff: ship the remaining bottom tray closeout parts to INFN.

10.Robert: get the corrected test reports and certs sent to INFN for the MCMs carried by Persis.


  1. MCM screening and inspection status. 
    1. SLAC looks on track to have 36 MCMs fully reworked and tested by Friday.  Only MCMs with connector height <= 3.91mm are being prepared.  SLAC will send Teledyne certs and corrected test data sheets for the MCMs that Persis hand carried.  SLAC will start preparations for reworking MCMs with gaps between connectors and boards.  It was decided that SLAC will ship as many MCMs as possible on Friday, as this is the only hope for not slowing down G&A.
    2. Mutsumi has improved the software to include the pitch-adapter and bias trace tests and combine all data onto one test sheet.  Since workers often forget to specify the correct MCM part number, it was suggested that the program should prompt the user for the part number on each MCM tested. 
  2. Dave Rich reported on the status of getting Teledyne production back online.  We have resolved procurement issues, and this morning SLAC received a revised proposal from Teledyne; this needs to go through SLAC system.  Dave estimated mid to late next week to get the PO done.  New pitch adapter designs will be tested early next week.  Before resuming production we will have a delta-MRR.  Dave intends to arrange a date today.  Richard Gobin and Diane Graeme will alternate as source inspectors 1 week at a time until a new full-time inspector is hired and trained.  Diamond Tool is quoting on the 2nd PA test fixture.  40 new storage cases and lids should be delivered by Monday, and a PR is in the works for 250 of them.
  3. Flex-circuit cable status.
    1. 75-ohm termination resistor rework is needed on the Tower-A cables.  Before we thought we could get by without the rework, by choosing for Tower-A only MCMs with large duty-cycle timing margin.  However, the other recent MCM problems have made that impractical.  Zentek is interested in doing the work, but they want us to push SLAC to complete payment for their previous work.
    2. Testing of cables for Tower-A.  Hiro tested them already, but they will have to be tested again after the rework.
    3. Status of Parlex issues: SLAC is holding daily meetings with Parlex. Lowell visited Tuesday, and Dave Nelson is taking over as the SLAC point of contact.  The barrel plating issue still has not been completely resolved.  Dave said that a chemical supplier is to visit Parlex to advise on etching.
    4. Second & third source status.  Titan is working on first articles of full-length cables, which are due December 1st.  A QA review of Titan will be carried out next week.  A PR for Cirrex is being reviewed and approved, for them to make prototypes of spliced cables.
  1. Status of other Tower-A deliverables from SLAC to INFN.
    1. Grid I/F hardware: studs are at Nyloc being reworked.  Arthur was there last week and will return next week.  Only 182 studs passed his thread go-no-go test.  Those are at Nyloc.  On the failed threads there are burrs that need to be cleaned, and some have too large pitch diameter and need to go back to AMT for rework.  Arthur will check them again before sending them back to AMT.
    2. Cable bending tools came in yesterday.  4 more sets are due by end of November.
    3. Steve: the inner shipping container, outer container, dummy load are nearly ready to ship.  He just needs screws to hold down the data logger.  Probably he can ship on Monday.  The assembly includes the cable holding plate, baseplate, and all hardware.  A new foam insert should be here today.  The original one was the wrong dimension.
    4. Peter: alignment tools and 10 stud sockets are in work.  They should be in on Monday from Diamond Tool. 
    5. Fasteners are coming in this week and next week for the top tray.  The dry fit check to the top tray in Pisa looks okay.
    6. Poron foam is not purchased yet.  Jeff will check up on it.
    7. Jeff will send a non-flight cable to Pisa next week for fit checks to the top tray.
    8. Jeff should receive flight-quality Kapton tape this week and will send it to Pisa.
    9. Jeff ordered 4-40 fasteners with locking mechanisms for the thermal straps.  The lead time is 3 weeks.  Can non-286 screws be used?  The EM vibe test used non A286 screws.  We could use Solithane for locking.
    10. All heat straps are at SLAC being QC’ed.  Also the bars.  The tower-A set plus some can be shipped next week.
    11. Solithane shipped out yesterday.  Jeff will send a pro-forma.  Probably it will arrive Sunday or Monday.
    12. New M2.5 100-degree torx+ screws should be in today; SLAC will test them and ship next week.  The old ones clearly had a fit problem.  Samples received of the new ones look and feel much better.
    13. The desiccant needs to be sent; also the ESD wrap.  A big shipment will be going out next week.
    14. SLAC needs to ship bias circuits tomorrow!  About 180 of them are ready.
    15. Tungsten was received in Italy from NASA, so that seems to be going well.
    16. 2 EGSE systems should be at SLAC ready to ship; Hiro needs to inspect them.  A 3rd is also owed to Italy but is being used right now for testing cables at SLAC.
    17. EGSE T/V cables: 1 complete set of long cables is in Alenia working.  They also have 3 half GASU cables.  Jeff needs to find and ship the other halves.  There should be 3 PDU cables in Italy.  Jeff and Pisa need to check on this.  According to the July 1 minutes 1 PDU cable was in transit at that time, and Jeff had 2 more to ship.
  2. Document approvals:
    1. Revised tray assembly drawings.  We need signatures from Darren and Jack, as well as all Tracker people.
    2. We need TD-03651 approval from Sandro.
    3. Sidewall procurement specs, PS-03467, needs approvals from Sandro, Riccardo and Ben Rodini.
  3. Additional BT closeouts and CC material: 6 + 4 bottom tray closeouts are ready to ship; 20 tray sets previously shipped to INFN.  One was used on the EM, and another was used on Tower-0.
  4. G&A status/progress:  They have put electronics on 4 trays.  The connector height matching is very tight, making the work slow. They are now proceeding with mounting ladders, and these trays will be ready for testing Monday.  INFN Pisa is working hard on adapting pockets for the MCMs on hand.  Nanda and Sandro estimated December 1 for the start of Tower-A assembly.
  5. Trays, bubbles, cuts, etc.  There will be an MRB tomorrow on Kapton bubbles and cutting.
  6. A.O.B.  The 2nd data logger is on its way back from Pisa to SLAC.  Sandro should have certificates for the 1st prepreg shipments in his email.  That prepreg is at Plyform.  The 2nd shipment needs to go out such that it arrives no later than a Wednesday in Pisa.